# DoubleTake: Geometry Guided Depth Estimation

### v1

@@ 3.5 Training data generation
A Ours without Hint MLP (as in SimpleRecon [54]) .0873 .0430 .0128 .1483 74.12 98.05
@@
E Ours w/ hint-based variable depth planes [72] .0821 .0405 .0121 .1432 76.67 98.19
F SimpleRecon [54] w/ TOCD [33] .0880 .0437 .0134 .1505 74.19 97.83
G Ours w/ single MLP for matching and hints .0773 .0371 .0112 .1381 79.56 98.35
H Ours (no hint) .0870 .0428 .0128 .1477 74.35 98.02
I Ours (incremental, fast) .0826 .0400 .0125 .1473 77.14 98.05
@@
Rendered depth from TSDF .2506 .1293 .3338 23.53 67.97
Densified rendered depth .1763 .0629 .2264 30.61 81.69
SimpleRecon [54] .1350 .0437 .1879 46.88 89.32
Ours (no hint) .1346 .0449 .1879 47.28 89.39
Ours (incremental) .1255 .0395 .1787 48.76 90.47
Ours (revisit) .1182 .0368 .1710 50.24 91.78
Ours (revisit, pose noise) .1199 .0372 .1725 49.46 91.56
@@
Table 3: Long-term hints using 3RScan. Ours (revisit) shows depth scores when we use the geometry from a previous visit to the same location as ‘hints’ for our current, incremental depth estimates. This mode beats previous baselines for this dataset, validating our proposal to retain long-term hints. Rendered depth uses the rendering of the prior TSDF as the current depth estimate. Densified rendered depth improves upon the above baseline with a network to fill in missing geometry. Full metrics and descriptions are in the supplementary. On the right we show we are robust to stale geometry: The chair has moved position since the hint TSDF was generated, but our system gracefully recovers.
Input GT Hint Confidence Ours SR [54]
@@
Figure 5: Qualitative results on 3RScan. Note the major improvement of our model over [54] on this challenging dataset, and how we can recover from misleading hints.
3.6 Implementation details
We train our model (and all our ablations) with a batch size of 16 on each of two Nvidia A100 GPUs. We use a learning rate of $41e{-}$ 4 that we drop by a factor of 10 at $70000$ and $80000$ steps. We train with an input dimension of $512\times 384$, and an output dimension of $256\times 192$. During training, we color augment RGB images. Our network follows that of [54], with the addition of our ‘Hint MLP’ as described above. This network uses an EfficientNetV2 S [64] encoder for the monocular image prior and the first two blocks of a ResNet18 [23] encoder for generating matching features for the cost volume. The decoder follows UNet++ [80]. Full architecture details are given in the supplementary material. Our training losses directly follow [54]. Our Hint MLP is small with two hidden layers, each comprising 12 neurons; this adds $2$ms to our runtime. We select source frames for the cost volume with the strategy and hyperparameters from [15]. For evaluation of online methods, we use online source views for each keyframe’s cost volume, so source frames come from the past in a sequence. For evaluation of offline methods, source frames may come from anywhere in a sequence. For both ‘ours’, SimpleRecon [54] and DeepVideoMVS [15] we use the same TSDF fusion method. We fuse depth maps to a maximum depth of $3.5$m into a TSDF volume of $2$cm resolution. For the FineRecon [62] baseline we use a $1$cm voxel resolution.
4 Evaluation
We evaluate with three challenging 3D datasets, all acquired with a handheld RGBD sensor. We train and evaluate on ScanNetV2 [12], which comprises 1,201 training scans, 312 validation scans, and 100 testing scans of indoor scenes. We additionally evaluate our ScanNetV2 models on 7-Scenes [57] without fine-tuning, following e.g. [54] and with the test split from [15]. We also evaluate on 3RScan [68]. This dataset captures the same environment in multiple separate scans, between which objects’ positions have changed. This tests our ability to use scans captured in the past as ‘hints’ for instantaneous depth estimates.
4.1 Mesh evaluation metrics
We follow existing works [2, 47] and report reconstruction metrics based on point to point distances on sampled point clouds. ScanNetV2 ground truth meshes are not complete, as the scan sequences don’t have full coverage. This means methods which overpredict geometry not in the ground truth get unfairly punished. TransformerFusion [2] use a visibility mask to trim predictions when computing prediction to ground-truth distances. However, these masks are over-sized, and include large areas of geometry which aren’t present in the ground truth mesh. We propose new visibility volumes using rendered depth maps of the ground-truth meshes, which much tighter fit the ground truth meshes. Figure 6 shows the difference between masks from [2] and our new proposed masks. More are shown in the supplementary. We show results on both sets of masks in Tables 5 and 4, and we will release our updated masks for reproducibility.
4.2 Depth and reconstruction performance
Depth estimation.
Table 1 shows our depth estimation performance on ScanNetV2 and 7Scenes in both incremental and offline modes (see Section 3.4 for details of these modes). Our incremental method outperforms all baseline methods across most scores. Our offline approach outperforms all competitors including our incremental method.
Reconstruction.
For meshing results we separate all methods into ‘online’ and ‘offline’: ‘Online’ methods can produce meshes instantaneously, while ‘offline’ methods are designed with the assumption that all frames are available at once. Tables 5 and 4 show reconstruction performance both ‘online’ and ‘offline’. Our incremental model sets a new state-of-the-art for online reconstruction performance, validating our approach to depth and reconstruction. See Figures 4 and 7 for qualitative comparisons with previous state-of-the-art. Ours (offline) obtains even better results; we present these in the second section of the table, where we compare against other offline reconstruction approaches e.g. [62, 61, 2]. Our approach is first or second best in most metrics. Figure 8 shows the benefit running ours offline can bring.
Figure 6: We introduce a more accurate mesh evaluation. (a) is the ground truth mesh, which contains many holes. (b) is an example predicted mesh, here from [62]. This is punished for being too complete, as [2]’s visibility mask (c) extends beyond the ground truth, giving high Acc error on their prediction (d). Our new masking (e) is tighter to the ground truth mesh, giving a more meaningful error (f).
[…]
Online DeepVideoMVS [15] No 6.49 6.97 6.73 .568 .595 .579
NeuralRecon [63] Yes 7.31 10.81 9.06 .453 .592 .511
SimpleRecon [54] (online) No 5.56 5.02 5.29 .631 .712 .668
Ours (incremental) No 4.92 5.49 5.20 .685 .701 .692
ATLAS [47] Yes 5.59 7.52 6.55 .671 .610 .637
Offline TransformerFusion [2] Yes 4.68 8.27 6.48 .698 .600 .644
VoRTX [61] Yes 4.38 7.23 5.80 .726 .651 .685
SimpleRecon [54] (offline) No 5.25 4.86 5.05 .654 .725 .687
FineRecon [62] Yes 4.92 5.06 4.99 .687 .737 .710
Ours (offline) No 4.15 4.85 4.50 .743 .734 .738
Table 4: Mesh Evaluation on ScanNetV2 with new visibility masks. We evaluate methods using the evaluation from [2] using visibility masks that more accurately represent the ground-truth mesh; see Sec. 4.1 and supplementary material for details. The ‘Volumetric’ category indicates whether a technique involves volumetric 3D reconstruction. Following [54], for other Multi-View Stereo (MVS) methods that generate solely depth maps, we applied conventional TSDF fusion for reconstruction. Note that VoRTX wins on accuracy, but its very sparse predictions give a poor completion score. Chamfer distance is the mean of accuracy and completion and F-Score is the harmonic mean of precision and recall. $\dagger$ Like Ours (offline), [54] (offline) uses source frames from both past and future.
[…]
Online RevisitingSI [25] No 14.29 16.19 15.24 .346 .293 .314
MVDepthNet [69] No 12.94 8.34 10.64 .443 .487 .460
GPMVS [24] No 12.90 8.02 10.46 .453 .510 .477
ESTDepth [38] No 12.71 7.54 10.12 .456 .542 .491
DPSNet [27] No 11.94 7.58 9.77 .474 .519 .492
DELTAS [59] No 11.95 7.46 9.71 .478 .533 .501
DeepVideoMVS [15] No 5.84 6.97 6.41 .639 .595 .615
NeuralRecon [63] Yes 5.09 9.13 7.11 .630 .612 .619
SimpleRecon [54] (online) No 5.72 5.02 5.37 .682 .712 .696
Ours (incremental) No 4.70 5.49 5.09 .730 .701 .714
Offline COLMAP [56, 55] No 10.22 11.88 11.05 .509 .474 .489
ATLAS [47] Yes 7.11 7.52 7.31 .679 .610 .640
3DVNet [52] Yes 6.73 7.72 7.22 .655 .596 .621
TransformerFusion [2] Yes 5.52 8.27 6.89 .729 .600 .655
VoRTX [61] Yes 4.31 7.23 5.77 .767 .651 .703
SimpleRecon [54] (offline)$\dagger$ No 5.37 4.86 5.12 .702 .725 .712
FineRecon [62] Yes 5.25 5.06 5.16 .779 .737 .756
Ours (offline) No 4.96 4.85 4.90 .752 .734 .742
Table 5: Mesh Evaluation on ScanNetV2. Here we use the evaluation and visibility masks from [2].
Long-term hints on 3RScan.
Our system can use long-term hints e.g. where we revisit an environment we have first observed at some time in the past. We use the 3RScan dataset [68] for this scenario, as this includes multiple scans of the same location at different times. We use the TSDF generated from a previous visit as the hint for the current, online depth estimates. See Table 3 for our results, where we see the benefit of using long-term hints vs baselines which don’t have access to hints, or an incremental version of our method. The table also includes figures showing how our method can gracefully cope with the situation where the scene has changed since the hint TSDF was generated.
@@ 4.2 Depth and reconstruction performance
NeuralRecon [63] SR [54] (online) Ours (incremental) Ground truth
@@
Figure 7: Qualitative results for meshing (incremental). Our method gives higher quality meshes than baselines e.g. [54] by using existing geometry for future predictions.
Timings and memory.
Our online, incremental system takes just 76.6ms to compute depth for a single frame, as measured on an Nvidia A100. The majority of this time (52.8ms) is running the forward pass of the depth network, while the remainder is generating the hint and updating the TSDF. A version of our model with smaller networks than [54] runs at 50.4ms per frame, see 2 in Table 2. This compares to e.g. 58ms per frame update of [54] or 90ms of [63]. Offline, we are faster than close competitor FineRecon [62], taking on average 13.8s per scene vs. [62]’s 48.1s. See Supplementary Material for a full breakdown of timings.
4.3 Ablations and variants
Table 2 shows ablations and variants of our approach in incremental mode. Row 2 doesn’t use a geometry hint or a Hint MLP at all in training or evaluation, so it is functionally equivalent to [54]. Ablations 2 and 2 replace our Hint MLP with alternatives methods; both of these score worse than ours (2). Our use of a separate Hint MLP has an additional advantage that we can cache the cost volume output for the second pass of offline mode. In 2 the network has no access to confidences, so it may incorrectly rely on under-construction subpar geometry. We then vary the format the geometry hint takes: ablation 2 forward-warps previous depth predictions to the current viewpoint. 2 and 2 are implementations in the spirit of previous publications [72, 33] which allow depth as input to an MVS system; see supplementary for full details. Finally, 2 is our full model but where we avoid giving a hint for any test-time frames; this performs similarly to 2, showing we are not disadvantaged in situations where no hint is available. Finally, Ours (2) outperforms all ablations.
Sensitivity to pose errors.
In the revisit scenario, to simulate errors in the relocalization algorithm, we add noise to the one-off alignment between the previous capture and the current capture. The result is shown in Table 3, and shows our tolerance to noisy alignments. See the supplementary for details.
Non-static scenes.
FineRecon [62] SR [54] (offline) Ours (offline) Ground truth
@@ 4.3 Ablations and variants
Figure 8: Qualitative results for meshing (offline, with two passes). Our method enables higher quality meshes than baselines e.g. [54] by running it twice.
@@ 4.4 Limitations
Like other methods which reconstruct 3D scenes by fusing depth maps e.g. [54], we only reconstruct geometry we have directly observed, and do not complete any hidden or out-of-frustum surfaces. Since we limit to $3.5$m the depths fused into the TSDF, our method has the most benefits when revisiting an area that has been previously observed close by. For example, if the camera is moving forward through a long corridor, our proposed geometric hints would only cover a small region of the image. Similarly to other multi-view stereo based methods, our method struggles with transparent and reflective surfaces.

### v2

@@ 4 Evaluation
DPSNet [im2019dpsnet] .1552 .0795 .0299 49.36 93.27 .1966 .1147 .0550 38.81 87.07
MVDepthNet [wang2018mvdepthnet] .1648 .0848 .0343 46.71 92.77 .2009 .1161 .0623 38.81 87.70
DELTAS [sinha2020deltas] .1497 .0786 .0276 48.64 93.78 .1915 .1140 .0490 36.36 88.13
GPMVS [hou2019multi] .1494 .0757 .0292 51.04 93.96 .1739 .1003 .0462 42.71 90.32
DeepVideoMVS, fusion [duzceker2021deepvideomvs]* .1186 .0583 .0190 60.20 96.76 .1448 .0828 .0335 47.96 93.79
SimpleRecon [sayed2022simplerecon] .0873 .0430 .0128 74.12 98.05 .1045 .0575 .0156 60.12 97.33
Table 1: Depth evaluation on ScanNetV2 and 7Scenes. Unless stated otherwise, predictions are computed incrementally, without access to future frames or frames from previous scans of the scene. We highlight the best, second-best and third-best methods per metric. Previous scores are from [duzceker2021deepvideomvs, sayed2022simplerecon]. *[duzceker2021deepvideomvs] was boosted by computing depths using three inference frames instead of two; they also use a custom 90/10 split.
4.1 Mesh evaluation metrics
We follow existing works [bozic2021transformerfusion, murez2020atlas] and report reconstruction metrics based on point to point distances on sampled point clouds. ScanNetV2 ground truth meshes are not complete, as the scan sequences don’t have full coverage. This means methods which overpredict geometry not in the ground truth get unfairly punished. TransformerFusion [bozic2021transformerfusion] use a visibility mask to trim predictions when computing prediction to ground-truth distances. However, these masks are over-sized, and include large areas of geometry which aren’t present in the ground truth mesh. We propose new visibility volumes using rendered depth maps of the ground-truth meshes, which much tighter fit the ground truth meshes. Figure 4 shows the difference between masks from [bozic2021transformerfusion] and our new proposed masks. More are shown in the supplementary. We show results on both sets of masks in Tables 2 and 4, and we will release our updated masks on acceptance for reproducibility.
4.2 Depth and reconstruction performance
Table 1 shows our depth estimation performance on ScanNetV2 and 7Scenes in both incremental and offline modes (see Section 3.4 for details of these modes). Our incremental method outperforms all baseline methods across most scores. Our offline approach, where we have previously seen the same sequence, outperforms all competitors including our incremental method. Tables 2 and 4 show reconstruction performance both in incremental and offline modes. Our incremental model sets a new state-of-the-art for online reconstruction performance, validating our approach to depth and reconstruction. See Figures 5 and 7 for qualitative comparisons with previous state-of-the-art. Ours (offline) obtains even better results; we present these in the second section of the table, where we compare against other offline reconstruction approaches e.g. [Stier_2023_ICCV, stier2021vortx, bozic2021transformerfusion]. Our approach is first or second best in most metrics. Figure 8 shows the benefit running ours offline can bring.
[…]
Online RevisitingSI [hu2019revisiting] No 14.29 16.19 15.24 .346 .293 .314
MVDepthNet [wang2018mvdepthnet] No 12.94 8.34 10.64 .443 .487 .460
GPMVS [hou2019multi] No 12.90 8.02 10.46 .453 .510 .477
ESTDepth [long2021multi] No 12.71 7.54 10.12 .456 .542 .491
DPSNet [im2019dpsnet] No 11.94 7.58 9.77 .474 .519 .492
DELTAS [sinha2020deltas] No 11.95 7.46 9.71 .478 .533 .501
DeepVideoMVS [duzceker2021deepvideomvs] No 5.84 6.97 6.41 .639 .595 .615
NeuralRecon [sun2021neuralrecon] Yes 5.09 9.13 7.11 .630 .612 .619
SimpleRecon [sayed2022simplerecon] (online) No 5.72 5.02 5.37 .682 .712 .696
Ours (incremental) No 4.70 5.49 5.09 .730 .701 .714
Offline COLMAP [schonberger2016pixelwise, schoenberger2016sfm] No 10.22 11.88 11.05 .509 .474 .489
ATLAS [murez2020atlas] Yes 7.11 7.52 7.31 .679 .610 .640
3DVNet [rich20213dvnet] Yes 6.73 7.72 7.22 .655 .596 .621
TransformerFusion [bozic2021transformerfusion] Yes 5.52 8.27 6.89 .729 .600 .655
VoRTX [stier2021vortx] Yes 4.31 7.23 5.77 .767 .651 .703
SimpleRecon [sayed2022simplerecon] ($\dagger$) No 5.37 4.86 5.12 .702 .725 .712
FineRecon [Stier_2023_ICCV] Yes 5.25 5.06 5.16 .779 .737 .756
Ours (offline) No 4.96 4.85 4.90 .752 .734 .742
Table 2: Mesh Evaluation on ScanNetV2. Here we use the evaluation from [bozic2021transformerfusion]. The ‘Volumetric’ category indicates whether a technique involves volumetric 3D reconstruction. Following [sayed2022simplerecon], for other Multi-View Stereo (MVS) methods that generate solely depth maps, we applied conventional TSDF fusion for reconstruction. Note that VoRTX wins on accuracy, but its very sparse predictions give a poor completion score. We have bolded the names of the most ‘telling’ metrics: Chamfer distance (the mean of accuracy and completion) and F-Score (harmonic mean of precision and recall). $\dagger$ Like Ours (offline), [sayed2022simplerecon] (offline) uses keyframes from anywhere in the sequence.
[…]
A Ours without hint MLP (as in [sayed2022simplerecon]) .0873 .0430 .0128 .1483 74.12 98.05
@@ 4.2 Depth and reconstruction performance
E Ours w/ cost volume modulation [poggi2022multi] .0855 .0427 .0131 .1481 75.07 97.84
F Ours w/ variable depth planes [Xin2023ISMAR] .0968 .0496 .0145 .1571 68.19 97.72
G SimpleRecon [sayed2022simplerecon] w/ TOCD [khan2023temporally] .0880 .0437 .0134 .1505 74.19 97.83
H Ours w/ single MLP for matching and hints .0773 .0371 .0112 .1381 79.56 98.35
I Ours (no hint) .0870 .0428 .0128 .1477 74.35 98.02
@@
Online DeepVideoMVS [duzceker2021deepvideomvs] No 6.49 6.97 6.73 .568 .595 .579
NeuralRecon [sun2021neuralrecon] Yes 7.31 10.81 9.06 .453 .592 .511
SimpleRecon [sayed2022simplerecon] (online) No 5.56 5.02 5.29 .631 .712 .668
Ours (incremental) No 4.92 5.49 5.20 .685 .701 .692
ATLAS [murez2020atlas] Yes 5.59 7.52 6.55 .671 .610 .637
Offline TransformerFusion [bozic2021transformerfusion] Yes 4.68 8.27 6.48 .698 .600 .644
VoRTX [stier2021vortx] Yes 4.38 7.23 5.80 .726 .651 .685
SimpleRecon [sayed2022simplerecon] (offline) No 5.25 4.86 5.05 .654 .725 .687
FineRecon [Stier_2023_ICCV] Yes 4.92 5.06 4.99 .687 .737 .710
Ours (offline) No 4.15 4.85 4.50 .743 .734 .738
Table 4: Mesh Evaluation on ScanNetV2 with new visibility masks. We evaluate methods using the evaluation from [bozic2021transformerfusion] using visibility masks that more accurately represent the ground-truth mesh; see Sec. 4.1 and supplementary material for details.
Input image Ground truth Ours SR [sayed2022simplerecon] DVMVS [duzceker2021deepvideomvs]
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Qualitative depth results on ScanNetV2. All methods are run incrementally, where we run at interactive speeds only with access to previous frames. We compare with [sayed2022simplerecon] and [duzceker2021deepvideomvs], the two closest-performing baselines. Our depth maps are more accurate with better small details (e.g. top) and overall geometry (e.g. bottom).
Long-term hints on 3RScan.
Our system can use long-term hints e.g. where we revisit an environment we have first observed at some time in the past. We use the 3RScan dataset [Wald2019RIO] for this scenario, as this includes multiple scans of the same location at different times. We use the TSDF generated from a previous visit as the hint for the current, online depth estimates. See Table 5 for our results, where we see the benefit of using long-term hints vs baselines which don’t have access to hints, or an incremental version of our method. The table also includes figures showing how our method can gracefully cope with the situation where the scene has changed since the hint TSDF was generated.
Timings and memory.
Our online, incremental system takes just 78.6ms to compute depth for a single frame, as measured on an Nvidia A100. The majority of this time (54.8ms) is running the forward pass of the depth network, while the remainder is generating the hint and updating the TSDF. This compares to e.g. 72ms per frame update of [sayed2022simplerecon] or 90ms of [sun2021neuralrecon]. Offline, we are faster than close competitor FineRecon [Stier_2023_ICCV], taking on average 13.8s per scene vs. [Stier_2023_ICCV]’s [TODO: ]s. See Supplementary Material for a full breakdown of timings.
[…]
Rendered depth .2506 .1293 .3338 23.53 67.97
SimpleRecon [sayed2022simplerecon] .1350 .0437 .1879 46.88 89.32
Densification .1763 .0629 .2264 30.61 81.69
Ours (no hint) .1339 .0441 .1874 47.71 89.44
Ours (incremental) .1240 .0394 .1776 49.82 90.63
Ours (revisit) .1169 .0369 .1700 51.33 91.87
Ours (revisit, pose noise) [TODO: ]
Ours (revisit, relocalization) [TODO: ]
@@
Table 5: Long-term hints using 3RScan. Ours (revisit) shows depth scores when we use the geometry from a previous visit to the same location as ‘hints’ for our current, incremental depth estimates. This mode beats previous baselines for this dataset, validating our proposal to retain long-term hints. Full metrics are in the supplementary. On the right we show we are robust to stale geometry: The chair has moved position since the hint TSDF was generated, but our system gracefully recovers.
Input GT Hint Confidence Ours SR [sayed2022simplerecon]
@@
Figure 6: Qualitative results on 3RScan. Note the major improvement of our model over [sayed2022simplerecon], and how we can recover from misleading hints.
4.3 Ablations and variants
Table 3 shows ablations and variants of our approach in incremental mode. Row 3 doesn’t use a geometry hint or a hint MLP at all in training or evaluation, so it is functionally equivalent to [sayed2022simplerecon]. Ablations 3 and 3 replace our Hint MLP with alternatives methods; both of these score worse than ours (3). Our use of a separate Hint MLP has an additional advantage that we can cache the cost volume output for the second pass of offline mode. In 3 the network has no access to confidences, so it may incorrectly rely on under-construction subpar geometry. We then vary the format the geometry hint takes: ablation 3 forward-warps previous depth predictions to the current viewpoint. 3, 3 and 3 are implementations in the spirit of previous publications [poggi2022multi, Xin2023ISMAR, khan2023temporally] which allow depth as input to an MVS system; see supplementary for full details. Finally, 3 is our full model but where we avoid giving a hint for any test-time frames; this performs similarly to 3, showing we are not disadvantaged in situations where no hint is available. Finally, Ours (3) outperforms all ablations.
Sensitivity to pose errors.
In the long-term hints case, to simulate errors in the relocalisation algorithm, we add noise to the one-off alignment between the previous view of the scene and the current view. In the incremental case, we add noise to every pose to simulate failures in the SLAM algorithm. [TODO: Include results]
Non-static scenes.
@@ 4.3 Ablations and variants
NeuralRecon [sun2021neuralrecon] SR [sayed2022simplerecon] (online) Ours (incremental) Ground truth
@@
Figure 7: Qualitative results for meshing (incremental). Our method gives higher quality meshes than baselines e.g. [sayed2022simplerecon] by using existing geometry for future predictions.
FineRecon [Stier_2023_ICCV] SR [sayed2022simplerecon] (offline) Ours (offline) Ground truth
@@
Figure 8: Qualitative results for meshing (offline, with two passes). Our method enables higher quality meshes than baselines e.g. [sayed2022simplerecon] by running it twice.
@@ 4.4 Limitations
Like other methods which reconstruct 3D scenes by fusing depth maps e.g. [sayed2022simplerecon], we only reconstruct geometry we have directly observed, and do not complete any hidden or out-of-frustum surfaces. Since we limit to $3$m the depths fused into the TSDF, our method has the most benefits when revisiting an area that has been previously observed close by. For example, if the camera is moving down a long corridor, our proposed geometric hints only cover a small subset of the visible area. Similarly to other multi-view stereo based methods, our method struggles with transparent and reflective surfaces.